OpenStudy (anonymous):

What s the simplified form of each expression? a^33/4 / a^8

OpenStudy (anonymous):

a^m/a^n=a^m-n so a^33/4-8 a^33-32/8 a^1/8

OpenStudy (anonymous):

hold on let me type my answer choices

OpenStudy (anonymous):

ok type

OpenStudy (anonymous):

A. A^1/4 B. A^66 C. 1/A^1/4 D. 1/ a ^66

OpenStudy (anonymous):

a^m/a^n =a^m-n so a^33/4-8 a^33-32/4 a^1/4 option a

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

giv me medal yar